Fermi National Laboratory
 

Operation Summary for February 26, 2010 - March 1, 2010

 

The Integrated Luminosity for the period from 2/22/10 to 3/1/10 was 62.27 pb-1.
NuMI reported receiving 7.24E18 protons on target during this same period.

Friday Feb. 26

The day shift began with Operations monitoring the Tevatron (TeV) store 7634 (88.96E30) and the Recycler stash (196.1E10), with the Antiproton Source off for maintenance work, with the Linac off for an RF station (LRF5) repair, with the Booster off and being prepared for access to work on RF stations, with the Main Injector operating normally, with the NuMI experiment being prepared for access, with the MiniBooNE experiment waiting for beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beam.

During the day shift, Booster and NuMI came out of access at 2:06 PM and Operations restored beam. Other areas of the accelerator complex performed minor maintenance work during these accesses.

During the evening shift, Operations terminated store 7634 at 6:17 PM. Operations prepared the CDF and D0 collision halls for access. The halls were out of access by 9:53 PM. Operations put the TeV through a wet squeeze at 11:08 PM.

Saturday Feb. 27

The midnight shift began with the TeV in a wet squeeze, with Operations monitoring the stash (399.4E10), with Pbar stacking (45.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TeV into HEP shot setup at 12:11 AM and established store 7638 at 1:29 AM with an Average Initial Luminosity (AIL) of 311.3E30. A P1 transfer line vacuum burst caused some minor downtime.

During the day shift, the I- Source (not being used) started to fail; operators contacted Linac technicians and managed to stabilize the source at a reduced output. The H- Source output started to drop as well; Operators tuned and raised its output.

During the evening shift, Operations terminated store 7638 at 4:52 PM and put the TeV into shot setup. Operations established store 7639 at 6:12 PM with an AIL of 316.9E30.

Sunday Feb. 28

The midnight shift began with Operations monitoring store 7639 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators tuned the H- Source and 750 KeV beam line to boost output. At 5:54 AM, Operations reported beam position problems in the Tevatron. Operators and TeV personnel investigated this problem through to the end of the shift.

During the day shift, operators continued to tune the H- Source. Operations terminated store 7639 at 10:19 AM and put the TeV into shot setup. Operations established store 7640 at 11:45 AM with an AIL of 310.4E30. A Linac Klystron RF station (KRF1) tripped off repeatedly due to a coil comparator power supply. Operators consulted with Linac personnel and got the station back online.

During the evening shift, LRF5 tripped off several times; it reset okay.

Monday March 1

The midnight shift began with Operations monitoring store 7640 and the stash, with Pbar stacking, with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7640 at 3:34 AM and put the TeV into shot setup. Operations established store 7641 at 5:10 AM with AIL of 297.9E30. Three ion pumps tripped off in the P1 transfer line and caused some minor downtime.

The Plans for Monday

The plans for today are to stack, stash, and store. Preaccelerator personnel will access the H- Source for 30 minutes of maintenance. CDF will access their collision hall in between stores.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/01/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory